OpenStudy (anonymous):

3 times the measure of an angle is 14 less than the measure of its complement. What is the measure of the angle? 19° 57° 71° 76°

OpenStudy (cwrw238):

first let the angle be x so how would you write the complement of x?

OpenStudy (anonymous):

complement means it equals 90 right?

OpenStudy (cwrw238):

complement of x = 90 - x

OpenStudy (anonymous):

so would i subtract 14 from 90 then divide by three?

OpenStudy (anonymous):

yes then u will get your answer

OpenStudy (anonymous):

but in doing that i get 25.3 repeating and that isnt an answer choice?

OpenStudy (cwrw238):

3x = (90 - x) - 14 3x = 76 - x 4x = 76 x = 19

OpenStudy (anonymous):

ohhhhhh thank you so much lo

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

x = (90 - x) - 14 3x = 76 - x 4x = 76 x = 19 this should help